A 500-N tightrope walker stands at the center of the rope such that each half of the rope makes an angle of 10 degrees with the horizontal.  What is the tension in the rope?
Answer #1
From the figure
    W = 2Tsinθ
therefore tension T = W/2sinθ = 500N / 2sin10= 1439.69 N